# Inverse problem for semidiscrete stochastic parabolic operators in arbitrary dimensions

### v1

@@ B.3. Returning to the original variable
$K(w)\triangleq$ $\sum_{i,j\in\llbracket 1,n\rrbracket}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1/2}\left(\mathcal{O}(h)+\mathcal{O}_{\lambda}((sh)^{2}+\mathcal{O}_{\lambda}(1)\right)|rD_{ij}^{2}w|^{2}\,dt$
$+\sum_{j\in\llbracket 1,n\rrbracket}\mathbb{E}\int_{Q_{j}^{\ast}}\left(s^{-1/2}\left(\mathcal{O}(h)+\mathcal{O}_{\lambda}((sh)^{2}+\mathcal{O}_{\lambda}(1)\right)+\mathcal{O}_{\lambda}(1)+s\mathcal{O}(sh)\right)|rD_{j}w|^{2}\,dt.$
Finally, there exist $\varepsilon_{3}(\lambda)>0$ and $s_{3}\geq 1$ such that combining (B.29) and (B.28) for $s\geq s_{3}$, $0<h<h_{3}$, and $0<sh\leq\varepsilon_{3}(\lambda)$, the desired inequality is obtained ∎
The two previous lemmas allow us to obtain a lower bound for the inequality (B.11) in terms of $w$. Now, we need to obtain an upper bound for the same inequality. To achieve this, we only need to estimate the second term on the right-hand side of the inequality (B.11), since the other terms are immediately available in terms of $w$.
Recalling that $w=\rho w$ and by Proposition 1.8 and Lemma A.1, we have $rD_{i}w=D_{i}z+\mathcal{O}_{\lambda}((sh)^{2})D_{i}z-\lambda s\varphi(\partial_{i}\psi)A_{i}(z)+s\mathcal{O}_{\lambda}((sh)^{2})A_{i}(z)$. We obtain the following
$\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}$ $s\lambda_{1}\varphi t_{r}^{i}(|D_{i}z|^{2})\,dt\leq\sum_{i=1}^{n}\left(\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\lambda_{1}\varphi t_{r}^{i}(|rD_{i}(w)|^{2})dt\right$.
$\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}$ $\int_{\Gamma^{+}}s\lambda_{1}\varphi t_{r}^{i}(|D_{i}z|^{2})\,dt\leq\sum_{i=1}^{n}\left(\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\lambda_{1}\varphi t_{r}^{i}(|rD_{i}(w)|^{2})dt\right$.
$\mathbb{E}\int_{Q}$ $|rf|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma^{+}}s\lambda_{1}t_{r}^{i}(|D_{i}z|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|z|^{2}\right|_{t=T}$
$\leq$ $C^{\prime\prime}_{s_{3},\epsilon_{3}}\left(\mathbb{E}\int_{Q}|rf|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\lambda_{1}\varphi t_{r}^{i}(|rD_{i}(w)|^{2})dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|rw|^{2}\right|_{t=T}\right)$
Consequently, applying the inequality mentioned above, (B.11), Lemma B.7 and Lemma B.8, we can determine that for $s_{0}=\max\{s_{1},s_{2},s_{3}\}$ and $h_{0}=\min\{h_{1},h_{2},h_{3}\}$, the desired Carleman estimate is obtained.
Appendix C Technical steps for the cross-product estimate
This section derives additional estimates not covered in EstimateCarlemanStochasticControl , which are essential to establish lemmas B.2 and B.3.
@@ Appendix C Technical steps for the cross-product estimate
$2\mathbb{E}\int_{Q}C_{6}zB_{1}z=-2\mathbb{E}\int_{Q}2s\Delta_{\gamma}(\varphi)zdz\triangleq\,I_{61}.$
@@
$X_{61}\triangleq$ $-2\left.\mathbb{E}\int_{Q}s(\lambda^{2}\varphi|\nabla_{\gamma}\psi|^{2}+\lambda\varphi\mathcal{O}(1))|z|^{2}\right|_{0}^{T}+\mathbb{E}\int_{Q}s\lambda\varphi\mathcal{O}(1)|dz|^{2}$
@@
Using the results of Appendix A and the boundary condition $z=0$ on $\partial_{i}Q$, we finally deduce

### v2

@@ A.3. Returning to the original variable
$K(w):=$ ${\sum_{i,j=1}^{n}}\mathbb{E}\int_{Q_{ij}^{\ast}}s^{-1/2}\left(\mathcal{O}(h)+\mathcal{O}_{\lambda}((sh)^{2}+\mathcal{O}_{\lambda}(1)\right)|rD_{ij}^{2}w|^{2}\,dt$
$+{\sum_{j=1}^{n}}\mathbb{E}\int_{Q_{j}^{\ast}}\left(s^{-1/2}\left(\mathcal{O}(h)+\mathcal{O}_{\lambda}((sh)^{2}+\mathcal{O}_{\lambda}(1)\right)+\mathcal{O}_{\lambda}(1)+s\mathcal{O}(sh)\right)|rD_{j}w|^{2}\,dt.$
Finally, there exist $\varepsilon_{3}(\lambda)>0$ and $s_{3}\geq 1$ such that combining (A.29) and (A.28) for $s\geq s_{3}$, $0<h<h_{3}$, and $0<sh\leq\varepsilon_{3}(\lambda)$, the desired inequality is obtained ∎
The two previous lemmas allow us to obtain a lower bound for the inequality (A.11) in terms of $w$. Now, we need to obtain an upper bound for the same inequality. To achieve this, we only need to estimate the second term on the right-hand side of the inequality (A.11), since the other terms are immediately available in terms of $w$.
From (Lecaros et al., 2026, Lemma 2.10), we first obtain that
$\begin{split}\mathbb{E}\int_{0}^{T}&\int_{G_{1}\cap\mathcal{M}_{i}^{\ast}}s\varphi\lambda^{2}|D_{i}z|^{2}\,dt\\ &\leq\,\mathcal{C}\left(\mathbb{E}\int_{0}^{T}\int_{G_{1}\cap\mathcal{M}_{i}^{\ast}}s\varphi\lambda^{2}\,|rD_{i}w|^{2}\,dt+\mathbb{E}\int_{0}^{T}\int_{G_{1}\cap\mathcal{M}_{i}^{\ast}}s\mathcal{O}_{\lambda}((sh)^{2})\,|D_{i}z|^{2}\,dt\right.\\ &\left.\mathbb{E}\int_{0}^{T}\int_{G_{1}\cap\mathcal{M}}s^{3}\varphi^{3}\lambda^{4}|z|^{2}\,dt+\mathbb{E}\int_{0}^{T}\int_{G_{1}\cap\mathcal{M}}s^{3}\mathcal{O}_{\lambda}((sh)^{2})\,|z|^{2}\,dt\right)\end{split}$
for $sh\leq\varepsilon_{0}$ and $i=1,..,n$. Moreover, taking into account that $\theta(t)$ satisfies (2.1), $\left.z\right|_{t=0}=0$ and using (Lecaros et al., 2026, Lemma 2.12), we have
$\begin{split}\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{G_{1}\cap\mathcal{M}_{i}^{\ast}}s\lambda^{2}_{1}\varphi e^{2s\varphi}\,|D_{i}w|^{2}\,dt\leq\sum_{i=1}^{n}&\mathbb{E}\int_{0}^{T}\int_{G_{0}\cap\mathcal{M}_{i}^{\ast}}s\lambda^{2}_{1}\varphi e^{2s\varphi}\,|D_{i}w|^{2}\,dt\\ \leq\mathcal{C}_{\tau_{3},\varepsilon_{0}}\left(\mathbb{E}\int_{0}^{T}\int_{G_{0}\cap\mathcal{M}}s^{3}\lambda^{2}\varphi^{3}e^{2s\varphi}\,|w|^{2}\,dt+\mathbb{E}\int_{0}^{T}\right.&\int_{G_{0}\cap\mathcal{M}}\lambda^{-2}e^{2s\varphi}|f|^{2}\\ &\left.+\left.\mathbb{E}\int_{G_{0}\cap\mathcal{M}}s\varphi e^{2s\varphi}\,|w|^{2}\right|_{t=T}\right).\end{split}$
Thus, combining the above two inequality, we can conclude that
(A.30) $\begin{split}\mathbb{E}\int_{0}^{T}&\int_{G_{1}\cap\mathcal{M}_{i}^{\ast}}s\varphi\lambda^{2}|D_{i}z|^{2}\,dt\leq\mathcal{C}_{\tau_{3},\varepsilon_{3}}\left(\mathbb{E}\int_{0}^{T}\int_{G_{0}\cap\mathcal{M}}s^{3}\lambda^{2}\varphi^{3}e^{2s\varphi}\,|w|^{2}\,dt\right.\\ &\left.+\mathbb{E}\int_{0}^{T}\int_{G_{0}\cap\mathcal{M}}\lambda^{-2}e^{2s\varphi}|f|^{2}+\left.\mathbb{E}\int_{G_{0}\cap\mathcal{M}}s\varphi e^{2s(T)\varphi}\,|w|^{2}\right|_{t=T}\right)\end{split}$
for $sh<\varepsilon_{0}$ and $i=1,\ldots,n$. Consequently, applying the inequality mentioned above, (A.11), Lemma A.7 and Lemma A.8, we can determine that for $s_{0}=\max\{s_{1},s_{2},s_{3}\}$ and $h_{0}=\min\{h_{1},h_{2},h_{3}\}$, the desired Carleman estimate is obtained.
Appendix B Technical steps for the cross-product estimate
This section derives additional estimates not covered in Lecaros et al. (2026), which are essential to establish lemmas A.2 and A.3.
@@ Appendix B Technical steps for the cross-product estimate
$2\mathbb{E}\int_{Q}C_{6}zB_{1}z=-2\mathbb{E}\int_{Q}2s\Delta_{\gamma}(\varphi)zdz:=\,I_{61}.$
@@
$X_{61}$= $-2\left.\mathbb{E}\int_{Q}s(\lambda^{2}\varphi|\nabla_{\gamma}\psi|^{2}+\lambda\varphi\mathcal{O}(1))|z|^{2}\right|_{0}^{T}+\mathbb{E}\int_{Q}s\lambda\varphi\mathcal{O}(1)|dz|^{2}$
@@
Using the results of Pérez (2024) and the boundary condition $z=0$ on $\partial_{i}Q$, we finally deduce
@@
Appendix C Technical steps for return to the original variable for the case boundary data in the right-hand side
Recalling that $w=\rho w$ and by Proposition 1.7 and (Boyer et al., 2010b, Lemma 3.7), we have $rD_{i}w=D_{i}z+\mathcal{O}_{\lambda}((sh)^{2})D_{i}z-\lambda s\varphi(\partial_{i}\psi)A_{i}(z)+s\mathcal{O}_{\lambda}((sh)^{2})A_{i}(z)$. We obtain the following
$\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}$ $s\lambda_{1}\varphi t_{r}^{i}(|D_{i}z|^{2})\,dt\leq\sum_{i=1}^{n}\left(\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\lambda_{1}\varphi t_{r}^{i}(|rD_{i}(w)|^{2})dt\right$.
$\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}$ $\int_{\Gamma}s\lambda_{1}\varphi t_{r}^{i}(|D_{i}z|^{2})\,dt\leq\sum_{i=1}^{n}\left(\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\lambda_{1}\varphi t_{r}^{i}(|rD_{i}(w)|^{2})dt\right$.
$\mathbb{E}\int_{Q}$ $|rf|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma}s\lambda_{1}t_{r}^{i}(|D_{i}z|^{2})\,dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|z|^{2}\right|_{t=T}$
$\leq$ $\mathcal{C}_{\tau_{3},\varepsilon_{3}}\left(\mathbb{E}\int_{Q}|rf|^{2}\,dt+\sum_{i=1}^{n}\mathbb{E}\int_{0}^{T}\int_{\Gamma_{+}}s\lambda_{1}\varphi t_{r}^{i}(|rD_{i}(w)|^{2})dt+\left.\mathbb{E}\int_{\mathcal{M}}s^{2}|rw|^{2}\right|_{t=T}\right).$